OpenStudy (anonymous):

Let A and B be events in a sample space S such that P(A)=.8 and P(B l A)=.6 Find P(A∩B).

OpenStudy (anonymous):

use the equation \[P(B|A) = \frac{ P(AnB) }{ P(A) }\]

OpenStudy (anonymous):

or the same equation written as \[P(A\cap B)=P(B|A)\times P(A)\]

OpenStudy (anonymous):

so .48?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

welcome